OpenStudy (anonymous):

the bases of a trapezoid are 22 and 12 respectively. The angles at the extremities of one base are 65 degrees and 40 degrees respectively. Find the Two legs. Anyone can help me using the law of sins

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

@.Sam.

OpenStudy (campbell_st):

ok... here is the solution, it uses similar triangles produce the sides you want to find, to form a triangle angle at the top is 75 by angle sum of a triangle  the 2 triangles are similar by angle angle test so \[\frac{22}{\sin(75)} = \frac{b}{\sin(65)}\] and then find a \[\frac{12}{\sin(75)} = \frac{a}{\sin(65)}\] the the side you need to find is b - a then the same process on the opposite side

OpenStudy (campbell_st):

oops forgot to say, the bases of the trapezoid are parallel... that's why the base angles are the same.

OpenStudy (anonymous):

it is the figure given?

OpenStudy (campbell_st):

well all you do is produce the sloping sides to form the triangle